# Flow Matching with Gaussian Process Priors for Probabilistic Time Series Forecasting

### v1

@@ Abstract
Recent advancements in generative modeling, particularly diffusion models, have opened new directions for time series modeling, achieving state-of-the-art performance in forecasting and synthesis. However, the reliance of diffusion-based models on a simple, fixed prior complicates the generative process since the data and prior distributions differ significantly. We introduce TSFlow, a conditional flow matching (CFM) model for time series that simplifies the generative problem by combining Gaussian processes, optimal transport paths, and data-dependent prior distributions. By incorporating (conditional) Gaussian processes, TSFlow aligns the prior distribution more closely with the temporal structure of the data, enhancing both unconditional and conditional generation. Furthermore, we propose conditional prior sampling to enable probabilistic forecasting with an unconditionally trained model. In our experimental evaluation on eight real-world datasets, we demonstrate the generative capabilities of TSFlow, producing high-quality unconditional samples. Finally, we show that both conditionally and unconditionally trained models achieve competitive results in forecasting benchmarks, surpassing other methods on 6 out of 8 datasets.
@@ 1 Introduction
In this work, we simplify the trajectories of generative time series modeling through informed priors, specifically using Gaussian Processes (GPs) as prior distributions within the CFM framework. By aligning the prior distribution more closely with the underlying temporal dynamics of real-world data, we simplify and shorten the probability paths. This approach not only reduces the complexity of the learned transformations but also enhances the model’s performance in conditional and unconditional generation tasks. Furthermore, we demonstrate how both conditionally trained and unconditionally trained models can be leveraged for probabilistic forecasting. We propose conditional prior sampling and demonstrate how to use guidance (Dhariwal & Nichol, 2021) to bridge the gap between conditional and unconditional generation, allowing for flexible application of unconditional models without the need for conditional training procedures.
@@
  • We introduce TSFlow, a novel time series generative model that incorporates conditional Gaussian Processes as informed prior distributions within the Conditional Flow Matching framework.
  • We demonstrate how these priors simplify the probability paths and improve the generative performance.
@@ 2 Background: Conditional Flow Matching
Conditional Flow Matching (CFM) was recently introduced as a framework for generative modeling by learning a flow field to transform one distribution into another (Lipman et al., 2022). The learned flow field then yields a transformation $\phi_{1}$ that maps a sample $\mathbf{x}_{0}\sim q_{0}$ from a prior distribution $q_{0}$, e.g., an isotropic Gaussian, to a transformed sample $\phi_{1}(\mathbf{x}_{0})$ that follows the data distribution $q_{1}$, i.e., $\phi_{1}(\mathbf{x}_{0})\sim q_{1}$. The flow field $\phi_{t}$ is parametrized through a vector field $u_{\bm{\theta}}$, which is learned in a simple regression task.
@@ 2.2 Couplings
$p_{t}(\mathbf{x}\mid\mathbf{x}_{0},\mathbf{x}_{1})=\mathcal{N}(t\mathbf{x}_{1}-(1-t)\mathbf{x}_{0},\sigma_{\mathrm{min}}^{2}\mathbf{I}),$ (4)
@@ 3 TSFlow
In this section, we present our main contributions: (1) TSFlow, a novel flow matching model tailored for probabilistic time series forecasting (see Fig. 1). TSFlow supports both unconditional and conditional generation, i.e., generation without or with partially observed time series. (2) We explore non-i.i.d. prior distributions within the CFM framework and demonstrate that incorporating a Gaussian process prior improves unconditional generation by simplifying the optimal transport problem. Further, (3) we show how an unconditional model can be employed for forecasting by conditioning it during inference via Langevin dynamics and guidance techniques. Finally, (4) we describe how to use data-dependent prior distributions in the form of Gaussian process regression to train TSFlow for conditional forecasting directly.
@@
Consider a univariate time series $\mathbf{y}\in\mathbb{R}^{L}$ of length $L$ from the data distribution $q_{1}(\mathbf{y})$. We will denote with $(\mathbf{y}^{p}\ \mathbf{y}^{f})=\mathbf{y}$ the split of the time series into an observed past $\mathbf{y}^{p}$ and an unknown future $\mathbf{y}^{f}$ we aim to predict.
@@
Then, we introduce a conditionally trained model that directly incorporates $\mathbf{y}^{p}$ to model the conditional distribution $p_{\bm{\theta}}(\mathbf{y}\mid\mathbf{y}^{p})$. To simplify notation, we will use $\mathbf{y}$ instead of $\mathbf{x}_{1}$. Furthermore, we denote the noisy time series as $\mathbf{x}_{0}$, and its (temporal) past and future $\mathbf{x}_{0}^{\mathbf{p}}$ and $\mathbf{x}_{0}^{\mathbf{f}}$, respectively.
@@ 3.1 Unconditional Modeling
To better capture the temporal correlations in the time series $\mathbf{y}$, we explore Gaussian process priors instead of the default isotropic Gaussian distribution. By aligning the prior distribution $q_{0}$ more closely with the data distribution $q_{1}$, we simplify the optimal transport problem and enhance the model’s ability to learn temporal patterns (Sec. 3.1.1). Then, we describe how to condition the unconditional model on observed data $\mathbf{y}^{p}$ during inference by incorporating past observations into the generation process via Langevin dynamics and guidance, effectively transforming it into a conditional one (Sec. 3.1.2 and 3.1.3).
@@ 3.1.1 Informed Prior Distributions
Previous work by Tong et al. (2023) have shown that pairing prior and data samples based on the optimal transport map accelerates training, reduces the number of Neural Function Evaluations (NFEs) required during inference, and enhances the model’s performance by shortening the conditional paths and straightening the learned velocity field.
We can further enhance this effect by choosing a domain-specific prior distribution $q_{0}$ closer to the data distribution $q_{1}$ than the standard isotropic Gaussian, yet similarly easy to specify and sample from. In the context of time series, we propose to employ Gaussian process (GP) priors $\mathcal{GP}(0,K)$ with a kernel function $K(t,t^{\prime})$ (Rasmussen & Williams, 2005). Gaussian processes are well-suited for modeling time series data because they naturally capture temporal correlations. By choosing $K$, we can incorporate dataset-specific structures and temporal patterns into the prior without requiring an extensive training process.
@@
$K_{\mathrm{SE}}(t,t^{\prime})=\exp\left(-\frac{d^{2}}{2\ell^{2}}\right),\ K_{\mathrm{OU}}(t,t^{\prime})=\exp\left(-\frac{|d|}{\ell}\right),\ \text{and}\ K_{\mathrm{PE}}(t,t^{\prime})=\exp\left(-\frac{2}{\ell^{2}}\sin^{2}(d)\right),$
where $d=t-t^{\prime}$ and $\ell$ is a non-negative parameter that adjusts the length scale of the kernel.
@@
Effect on the Optimal Transport Problem.
Our results in Fig. 2 show that Gaussian process priors reduce the Wasserstein distance between the prior and data distribution if the kernel exploits characteristics of the data. When we parametrize the prior with SE and OU kernels, the Wasserstein distance to the data distribution decreases substantially, although it slowly converges to an isotropic prior as the sequence length increases. Notably, the periodic kernel achieves the largest reduction in distance on these datasets with only marginal growth as the sequence length increases because it effectively captures the periodicity — a common feature of real-world time series — especially well. These findings support our hypothesis that an informed choice of prior distribution can substantially shorten the probability paths and directly affect the difficulty of the learning problem.
@@ 3.1.2 Conditional Prior Sampling
To adapt the unconditional model for conditional generation after training, we propose conditional prior sampling, which conditions the prior distribution as $q_{0}(\mathbf{x}_{0}\mid\mathbf{y}^{p})$. It is important to note that this only conditions the prior distribution, not the generation process itself, which we discuss in Sec. 3.1.3.
@@
$\mathbf{x}_{0}^{(i+1)}=\mathbf{x}_{0}^{(i)}-\tau\nabla_{\mathbf{x}_{0}}\log q_{0}(\mathbf{x}_{0}^{(i)}\mid\mathbf{y}^{p})+\sqrt{2\tau}\xi_{i}\text{ with }\xi_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (7)
where $\tau$ is a fixed step size. With sufficient iterations, Eq. 7 converges to the conditional distribution (see Durmus & Moulines (2017) for a convergence analysis).
@@ 3.1.3 Guided Generation
To perform conditional generation more effectively, we can employ guidance techniques (Dhariwal & Nichol, 2021) to modify the score function of the model, directly conditioning the generation process on the observed data $\mathbf{y}^{\mathbf{p}}$. Unlike the previous method, where we only conditioned the prior distribution, we now adjust the dynamics of the generation process itself.
@@
To incorporate this guidance into the generation process, we adjust the vector field $u_{\bm{\theta}}$ by subtracting the guidance score, scaled by a factor $s$:
$\tilde{u}_{\bm{\theta}}(t,\mathbf{x}_{t})=u_{\bm{\theta}}(t,\mathbf{x}_{t})-s\cdot\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}^{p}\mid\mathbf{x}_{t}).$ (11)
We provide a detailed derivation of $\tilde{u}_{\bm{\theta}}(\mathbf{x}_{t},t)$ in App. C. By integrating this modified vector field $\tilde{u}_{\bm{\theta}}$ over time, we generate samples that are conditioned on the observed data $\mathbf{y}^{\mathbf{p}}$. The parameter $s$ allows us to control the influence of the conditioning, with larger values of $s$ resulting in samples that more closely resemble the observed past. To model the guidance score, we again use an asymmetric Laplace distribution centered around the output of the flow as before:
@@ 3.2 Conditional Modeling
TSFlow imposes additional structure on the prior distribution $q_{0}$ in the form of a Gaussian process, a non-parametric time series model. We build upon this and propose an alternative training and inference approach for conditional generation, i.e., forecasting, with TSFlow. Instead of only conditioning the sampling process on an observed sample $\mathbf{y}^{p}$, we additionally condition the prior distribution on the observed past data by approximating $q_{0}(\mathbf{x}_{0}\mid\mathbf{y}^{p})$ with $q_{0}(\mathbf{x}_{0}\mid\mathbf{y}^{p})$.
@@
which follows by assuming independence of $\mathbf{x}^{p}$ and $\mathbf{x}^{f}$ given $\mathbf{y}^{p}$. We choose $q_{0}(\mathbf{x}^{p}_{0}\mid\mathbf{y}^{p})=\mathcal{N}(\mathbf{x}^{p}_{0}\mid\mathbf{y}^{p},{\bm{I}})$ to retain the information from the observation. For the unobserved future part, we employ GPR to model the conditional distribution (Rasmussen & Williams, 2005). Specifically,
@@ 4 Experiments
In this section, we present our empirical results and compare TSFlow against various baselines using real-world datasets. Our primary objectives are threefold: First, to determine whether the generative capabilities are competitive to other generative frameworks. Second, to investigate the impact of different prior distributions on the results. Third, to evaluate how TSFlow compares to other models in probabilistic forecasting. Additionally, we aim to explore how conditional prior sampling of the unconditional version of TSFlow fares against its conditionally trained counterpart.
@@
To assess the generative capabilities of TSFlow in an unconditional setting, we calculate the 2-Wasserstein distance between the synthetic and real samples. Specifically, we generate 10,000 samples using our defined context length for this evaluation. Additionally, to compare the quality of the synthetic samples in a downstream task, we employ the Linear Predictive Score (LPS) (Kollovieh et al., 2023), which is defined as the (real) test CRPS of a linear regression model trained on synthetic samples.
@@ 4.1.1 Generative Capabilities
TSDiff 100 $3.112$±$\pm 0.32$ $0.085$±$\pm 0.129$ $29.556$±$\pm 1.445$ $6.095$±$\pm 0.292$ $5.422$±$\pm 1.173$ $7.228$±$\pm 0.128$ $67.281$±$\pm 6.992$ $183.833$±$\pm 7.741$
TSFlow Isotropic $4$ $2.931$±$\pm 0.246$ $\underline{0.021}$±$\pm 0.010$ $27.320$±$\pm 0.910$ $6.745$±$\pm 0.360$ $5.059$±$\pm 0.376$ $7.616$±$\pm 0.146$ $63.576$±$\pm 0.634$ $193.809$±$\pm 2.701$
OU $4$ $2.588$±$\pm 0.078$ $0.037$±$\pm 0.018$ $\mathbf{24.133}$±$\pm 0.648$ $6.203$±$\pm 0.183$ $4.389$±$\pm 0.161$ $7.292$±$\pm 0.232$ $62.038$±$\pm 0.897$ $\pm 1.913$
SE $4$ $2.540$±$\pm 0.090$ $0.035$±$\pm 0.016$ $\underline{24.455}$±$\pm 0.954$ $6.351$±$\pm 0.220$ $\mathbf{4.346}$±$\pm 0.126$ $7.300$±$\pm 0.224$ $\mathbf{61.763}$±$\pm 0.511$ $190.359$±$\pm 1.511$
PE $4$ $\mathbf{2.501}$±$\pm 0.091$ $0.033$±$\pm 0.012$ $25.712$±$\pm 0.437$ $6.353$±$\pm 0.184$ $4.456$±$\pm 0.140$ $7.411$±$\pm 0.210$ $\underline{61.912}$±$\pm 0.744$ $194.178$±$\pm 0.938$
TSFlow Isotropic $16$ $2.729$±$\pm 0.218$ $\mathbf{0.018}$±$\pm 0.008$ $27.235$±$\pm 0.911$ $6.078$±$\pm 0.171$ $4.584$±$\pm 0.106$ $7.403$±$\pm 0.205$ $62.990$±$\pm 0.935$ $187.921$±$\pm 2.688$
OU $16$ $2.659$±$\pm 0.130$ $0.029$±$\pm 0.015$ $25.999$±$\pm 1.352$ $\mathbf{5.930}$±$\pm 0.124$ $4.392$±$\pm 0.062$ $\underline{7.177}$±$\pm 0.159$ $64.800$±$\pm 2.048$ $\underline{179.398}$±$\pm 4.839$
SE $16$ $2.615$±$\pm 0.100$ $0.029$±$\pm 0.015$ $26.001$±$\pm 1.814$ $\underline{5.967}$±$\pm 0.127$ $\underline{4.369}$±$\pm 0.064$ $\mathbf{7.147}$±$\pm 0.144$ $63.860$±$\pm 2.705$ $\mathbf{176.837}$±$\pm 9.782$
PE $16$ $\underline{2.532}$±$\pm 0.056$ $0.027$±$\pm 0.010$ $26.501$±$\pm 0.920$ $6.158$±$\pm 0.181$ $4.504$±$\pm 0.091$ $7.202$±$\pm 0.097$ $63.775$±$\pm 3.574$ $186.977$±$\pm 1.469$

### v2

@@ Abstract
Recent advancements in generative modeling, particularly diffusion models, have opened new directions for time series modeling, achieving state-of-the-art performance in forecasting and synthesis. However, the reliance of diffusion-based models on a simple, fixed prior complicates the generative process since the data and prior distributions differ significantly. We introduce TSFlow, a conditional flow matching (CFM) model for time series combining Gaussian processes, optimal transport paths, and data-dependent prior distributions. By incorporating (conditional) Gaussian processes, TSFlow aligns the prior distribution more closely with the temporal structure of the data, enhancing both unconditional and conditional generation. Furthermore, we propose conditional prior sampling to enable probabilistic forecasting with an unconditionally trained model. In our experimental evaluation on eight real-world datasets, we demonstrate the generative capabilities of TSFlow, producing high-quality unconditional samples. Finally, we show that both conditionally and unconditionally trained models achieve competitive results across multiple forecasting benchmarks.
@@ 1 Introduction
In this work, we simplify the trajectories of generative time series modeling through informed priors, specifically using Gaussian Processes (GPs) as prior distributions within the CFM framework. We align the prior distributions with the underlying temporal dynamics of real-world data via optimal transport couplings. This approach not only reduces the complexity of the learned transformations but also enhances the model’s performance in conditional and unconditional generation tasks. Furthermore, we demonstrate how both conditionally and unconditionally trained models can be leveraged for probabilistic forecasting. We propose conditional prior sampling and demonstrate how to use guidance (Dhariwal & Nichol, 2021) to bridge the gap between conditional and unconditional generation, allowing for flexible application of unconditional models without the need for specialized conditional training procedures.
@@
  • We introduce TSFlow, a novel generative model for time series that incorporates conditional Gaussian Processes as informed prior distributions within the Conditional Flow Matching framework.
  • We demonstrate how these priors align with the data distribution and improve the unconditional generative performance.
@@ 2 Background: Conditional Flow Matching
Conditional Flow Matching (CFM) was recently introduced as a framework for generative modeling by learning a flow field to transform one distribution into another (Lipman et al., 2022). The learned flow field then yields a transformation $\phi_{1}$ that maps a sample $\mathbf{x}_{0}\sim q_{0}$ from a prior distribution $q_{0}$, e.g., an isotropic Gaussian, to a transformed sample $\phi_{1}(\mathbf{x}_{0})$ that follows the data distribution $q_{1}$, i.e., $\phi_{1}(\mathbf{x}_{0})\sim q_{1}$. The flow field $\phi_{t}$ is parametrized through a vector field $u_{\bm{\theta}}$, which is learned in a simple regression task. In the following, we give a short exposition of flow matching. For a thorough introduction to flow matching and diffusion, we refer the reader to Luo (2022); Lipman et al. (2022).
@@ 2.2 Couplings
$p_{t}(\mathbf{x}\mid\mathbf{x}_{0},\mathbf{x}_{1})=\mathcal{N}(t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0},\sigma_{\mathrm{min}}^{2}\mathbf{I}),$ (4)
@@ 3 TSFlow
In this section, we present our main contributions: (1) TSFlow, a novel flow matching model tailored for probabilistic time series forecasting (see Fig. 1). TSFlow supports both unconditional and conditional generation, i.e., generation without or with partially observed time series. (2) We explore non-i.i.d. prior distributions within the CFM framework and demonstrate that incorporating a Gaussian process improves unconditional generation by aligning the prior with the data distribution. Further, (3) we show how an unconditional model can be employed for forecasting by conditioning it during inference via Langevin dynamics and guidance techniques. Finally, (4) we describe how to use data-dependent prior distributions in the form of Gaussian process regression to train TSFlow for conditional forecasting directly.
@@
Consider a regularly sampled univariate time series $\mathbf{y}\in\mathbb{R}^{L}$ of length $L$ drawn from the data distribution $q_{1}(\mathbf{y})$ (corresponding to dimension $d$, as detailed in Sec. 2). We denote the time series as $\mathbf{y}=(\mathbf{y}^{p}\ \mathbf{y}^{f})$, where $\mathbf{y}^{p}$ represents the observed past and $\mathbf{y}^{f}$ the unknown future we aim to predict. To avoid ambiguity with the flow-matching time parameter $t$, we introduce $\tau$ as the time index for positions within the time series data. Here, the flow-matching time $t$ parameterizes the progression along the transformation path from the prior to the target time series distribution, while $\tau$ is orthogonal and refers to specific time points in the observed data.
@@
Then, we introduce a conditionally trained model that directly incorporates $\mathbf{y}^{p}$ to model the conditional distribution $p_{\bm{\theta}}(\mathbf{y}\mid\mathbf{y}^{p})$. To simplify notation, we denote $\mathbf{x}_{1}$ as $\mathbf{y}$ in the following. Furthermore, we denote the noisy time series as $\mathbf{x}_{0}$, and its (temporal) past and future $\mathbf{x}_{0}^{p}$ and $\mathbf{x}_{0}^{f}$, respectively.
@@ 3.1 Unconditional Modeling
Algorithm 1 Unconditional Training of TSFlow
1:Input: Prior distribution $q_{0}$, data distribution $q_{1}$, noise levels $\sigma_{t}$ network $u_{\bm{\theta}}$
2:for $\mathrm{iteration}=1,\dots$ do
5:     $t\sim\mathcal{U}(0,1)$    $\triangleright$ sample random time step $t$
7:     $\mathbf{x}_{t}\sim\mathcal{N}(\bm{\mu}_{t},\sigma_{\mathrm{t}}^{2}\mathbf{I})$triangleright sample from $p_{t}(\cdot\mid\mathbf{z})$
8:     $\mathcal{L}({\bm{\theta}})\leftarrow\|u_{\bm{\theta}}(t,\mathbf{x}_{t})-u_{t}(\mathbf{x}_{t}\mid\mathbf{z})\|^{2}$triangleright regress conditional vector field
10:end for
11:Return: $u_{\bm{\theta}}$
@@
To better capture the temporal correlations in the time series $\mathbf{y}$, we explore Gaussian process priors instead of the default isotropic Gaussian distribution. By introducing temporal structure and aligning the prior distribution $q_{0}$ more closely with the data distribution $q_{1}$ via optimal transport couplings, we enhance the model’s ability to learn temporal patterns (Sec. 3.1.1). Then, we describe how to condition the unconditional model on observed data $\mathbf{y}^{p}$ during inference by incorporating past observations into the generation process via Langevin dynamics and guidance, effectively transforming it into a conditional one (Sec. 3.1.2 and 3.1.3).
@@ 3.1.1 Informed Prior Distributions
Previous work by Tong et al. (2023) has shown that pairing prior and data samples based on the optimal transport map accelerates training, reduces the number of Neural Function Evaluations (NFEs) required during inference, and enhances the model’s performance by shortening the conditional paths and straightening the learned velocity field.
We can further enhance this effect by choosing a domain-specific prior distribution $q_{0}$ closer to the data distribution $q_{1}$ than the standard isotropic Gaussian, yet similarly easy to specify and sample from. In the context of time series, we propose to employ Gaussian process (GP) priors $\mathcal{GP}(0,K)$ with a kernel function $K(\tau,\tau^{\prime})$ (Rasmussen & Williams, 2005). Gaussian processes are well-suited for modeling time series data because they naturally capture temporal correlations. By choosing $K$, we can incorporate dataset-specific structures and temporal patterns into the prior without requiring an extensive training process.
@@
$K_{\mathrm{SE}}(\tau,\tau^{\prime})=\exp\Bigl{(}-\frac{d^{2}}{2\ell^{2}}\Bigr{)},\ K_{\mathrm{OU}}(\tau,\tau^{\prime})=\exp\Bigl{(}-\frac{|d|}{\ell}\Bigr{)},\ \text{and }K_{\mathrm{PE}}(\tau,\tau^{\prime})=\exp\Bigl{(}-\frac{2}{\ell^{2}}\sin^{2}(d)\Bigr{)},$
where $d=\tau-\tau^{\prime}$ and $\ell$ is a non-negative parameter that adjusts the length scale of the kernel.
@@ 3.1.2 Conditional Prior Sampling
To adapt the unconditional model for conditional generation post-training, we propose conditional prior sampling, which conditions the prior distribution as $q_{0}(\mathbf{x}_{0}\mid\mathbf{y}^{p})$. Given a sample $\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}\mid\mathbf{y}^{p})$, we use it as the initial condition for our vector field to generate new samples from the distribution $\mathbf{y}\mid\mathbf{x}_{0}$. It is important to note that this only conditions the prior distribution, not the generation process itself, which we discuss in Sec. 3.1.3.
@@
$\mathbf{x}_{0}^{(i+1)}=\mathbf{x}_{0}^{(i)}-\eta\nabla_{\mathbf{x}_{0}}\log q_{0}(\mathbf{x}_{0}^{(i)}\mid\mathbf{y}^{p})+\sqrt{2\eta}\xi_{i}\text{ with }\xi_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (7)
where $\eta$ is a fixed step size. With sufficient iterations, Eq. 7 converges to the conditional distribution (see Durmus & Moulines (2017) for a convergence analysis).
@@ 3.1.3 Guided Generation
To perform conditional generation more effectively, we can employ guidance techniques (Dhariwal & Nichol, 2021; Kollovieh et al., 2023) to modify the score function of the model, directly conditioning the generation process on the observed data $\mathbf{y}^{p}$. Unlike the previous method, where we only conditioned the prior distribution, we now adjust the dynamics of the generation process itself.
@@
To incorporate this guidance into the generation process, we adjust the vector field $u_{\bm{\theta}}$ and subtract the guidance score, scaled by a factor $s$:
$\tilde{u}_{\bm{\theta}}(t,\mathbf{x}_{t})=u_{\bm{\theta}}(t,\mathbf{x}_{t})-s\sigma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}^{p}\mid\mathbf{x}_{t}).$ (11)
We provide a detailed derivation of $\tilde{u}_{\bm{\theta}}(\mathbf{x}_{t},t)$ in App. C. By integrating this modified vector field $\tilde{u}_{\bm{\theta}}$ over time, we generate samples that are conditioned on the observed data $\mathbf{y}^{p}$. The parameter $s$ allows us to control the influence of the conditioning, with larger values of $s$ resulting in samples that more closely resemble the observed past. To model the guidance score, we again follow (Kollovieh et al., 2023) and use an asymmetric Laplace distribution centered around the output of the flow as before:
@@ 3.2 Conditional Modeling
TSFlow imposes additional structure on the prior distribution $q_{0}$ in the form of a Gaussian process, a non-parametric time series model. We build upon this and propose an alternative training and inference approach for conditional generation, i.e., forecasting, with TSFlow. Instead of only conditioning the sampling process on an observed sample $\mathbf{y}^{p}$, we additionally condition the prior distribution on the observed past data by factorizing $q_{0}(\mathbf{x}_{0}\mid\mathbf{y}^{p})$ with $q_{0}(\mathbf{x}^{p}_{0}\mid\mathbf{y}^{p})\,q_{0}(\mathbf{x}^{f}_{0}\mid\mathbf{y}^{p})$.
@@
which follows by assuming independence of $\mathbf{x}^{p}$ and $\mathbf{x}^{f}$ given $\mathbf{y}^{p}$. We choose $q_{0}(\mathbf{x}^{p}_{0}\mid\mathbf{y}^{p})=\delta(\mathbf{x}^{p}_{0}-\mathbf{y}^{p})$ to retain the information from the observation. For the unobserved future part, we employ GPR to model the conditional distribution (Rasmussen & Williams, 2005). Specifically,
@@ 4 Experiments
In this section, we present our empirical results and compare TSFlow against various baselines using real-world datasets. Our primary objectives are threefold: First, to determine whether the generative capabilities are competitive to other generative frameworks. Second, to investigate the impact of different prior distributions on the performance. Third, to evaluate how TSFlow compares to other models in probabilistic forecasting. Additionally, we aim to explore how conditional prior sampling of the unconditional version of TSFlow fares against its conditionally trained counterpart.
@@
To assess the generative capabilities of TSFlow in an unconditional setting, we calculate the 2-Wasserstein distance between the synthetic and real samples. Specifically, we generate 10,000 samples using our defined context length for this evaluation. Additionally, to compare the quality of the synthetic samples in a downstream task, we employ the Linear Predictive Score (LPS) (Kollovieh et al., 2023), which is defined as the (real) test CRPS of a linear regression model trained on synthetic samples (see Sec. A.5 for more details).
@@ 4.1.1 Generative Capabilities
TSDiff $100$ $3.112$±$\pm 0.32$ $0.057$±$\pm 0.035$ $29.556$±$\pm 1.445$ $6.095$±$\pm 0.292$ $5.422$±$\pm 1.173$ $7.228$±$\pm 0.128$ $67.281$±$\pm 6.992$ $214.981$±$\pm 20.16$
TSFlow Isotropic $4$ $2.929$±$\pm 0.051$ $0.027$±$\pm 0.003$ $27.264$±$\pm 0.232$ $6.539$±$\pm 0.073$ $5.193$±$\pm 0.097$ $7.515$±$\pm 0.081$ $62.019$±$\pm 0.587$ $211.369$±$\pm 1.545$
OU $4$ $2.696$±$\pm 0.068$ $0.026$±$\pm 0.004$ $\mathbf{23.762}$±$\pm 0.113$ $6.509$±$\pm 0.051$ $4.564$±$\pm 0.075$ $7.283$±$\pm 0.098$ $\mathbf{60.338}$±$\pm 0.075$ $203.970$±$\pm 7.638$
SE $4$ $2.704$±$\pm 0.056$ $0.027$±$\pm 0.003$ $\underline{23.786}$±$\pm 0.041$ $6.543$±$\pm 0.049$ $\underline{4.388}$±$\pm 0.016$ $7.375$±$\pm 0.082$ $\underline{60.452}$±$\pm 0.388$ $205.748$±$\pm 4.614$
PE $4$ $2.695$±$\pm 0.033$ $0.027$±$\pm 0.003$ $27.029$±$\pm 0.239$ $6.704$±$\pm 0.070$ $4.492$±$\pm 0.045$ $7.450$±$\pm 0.021$ $60.951$±$\pm 0.373$ $205.856$±$\pm 4.834$
TSFlow Isotropic $16$ $2.752$±$\pm 0.024$ $\underline{0.025}$±$\pm 0.003$ $26.332$±$\pm 0.223$ $\underline{5.910}$±$\pm 0.052$ $4.495$±$\pm 0.055$ $7.166$±$\pm 0.086$ $65.939$±$\pm 0.833$ $204.269$±$\pm 1.712$
OU $16$ $2.635$±$\pm 0.045$ $\mathbf{0.023}$±$\pm 0.003$ $25.222$±$\pm 0.259$ $\mathbf{5.893}$±$\pm 0.083$ $4.402$±$\pm 0.065$ $\mathbf{6.990}$±$\pm 0.039$ $64.871$±$\pm 1.083$ $\underline{200.808}$±$\pm 6.350$
SE $16$ $\underline{2.625}$±$\pm 0.052$ $\mathbf{0.023}$±$\pm 0.003$ $25.223$±$\pm 0.340$ $5.937$±$\pm 0.064$ $4.429$±$\pm 0.064$ $\underline{6.991}$±$\pm 0.054$ $64.172$±$\pm 0.690$ $\mathbf{199.880}$±$\pm 7.743$
PE $16$ $\mathbf{2.542}$±$\pm 0.049$ $\mathbf{0.023}$±$\pm 0.003$ $25.472$±$\pm 0.102$ $5.992$±$\pm 0.061$ $\mathbf{4.362}$±$\pm 0.032$ $7.051$±$\pm 0.065$ $64.102$±$\pm 0.723$ $200.829$±$\pm 8.673$